1) A plug flow reactor has a flow volume of 20%. If the initial concentration was 700ppm and we are dropping to 40ppm. calculate the time required to reach the 40ppm. Note that the tank size is 10 tonnes and the flow rate is 1 tonne per minute.   

2) Calculate the same above with a mixed volume of 80%
Answer #1

1.

The plug flow tank size is 10 ton

therefore total volume available for flow (V) = 0.2*10 = 2 ton

The flow rate (F) =1 ton/min

Residence time (t) = V/F = 2/1

t = 2 min

In the tank, 2 min are taken to drop the concentration from 700 ppm to 40 ppm.

2.  

The mixed volume is V = 0.8*10 = 8 ton

flow rate (F) =1 ton/min

Residence time (t) = V/F = 8/1

t = 8 min